At one corner of a 3cm by 4cm rectangle is placed a charge of -20pc and at the two adjacent corner are charge of 10pc .calculate the potential at the fourth corner.


1
Expert's answer
2022-02-02T12:13:49-0500







Potential due to point charge

"V=k_c\frac{q}{r}"


Where "V=" Electric potential

"q=" point charge

"r=" Distance between any point around the charge to the point charge

"k_c=" Coulombs constant "k_c=9\u00d710^9N"



Potential due to "+10pc" at A


"V_{+10A}=\frac{9×10^9×10×10^{-12}}{3×10^{-2}}"

"=3V"


Potential due to +10pc at C


"V_{+10C}=\frac{9×10^9×10×10^{-12}}{4×10^{-2}}=2.25V"

PB"=\sqrt{3^2+4^2}=5"


Potential due to -20pc at B


"V_{-20B}=\frac{-9×10^9×20×10^{-12}}{5×10^{-2}}" "=-3.6V"



Potential at P"=3+2.25-3.6"

"=1.65V"